OpenStudy (anonymous):

how do you find the minimum and maximum points of xe^-2x?

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

did you get the derivative yet?

OpenStudy (anonymous):

yeoo

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by "what" i mean "what did you get?"

OpenStudy (anonymous):

-2xe^-2x + e^-2x

OpenStudy (anonymous):

ok good, now factor out the common factor of \(e^{-2x}\)

OpenStudy (anonymous):

so x=-1/2

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

coool thanks

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[x=\frac{1}{2}\] is what you get when you solve

OpenStudy (anonymous):

that will give you the local max

OpenStudy (anonymous):

\[(1-2x)e^{-2x}=0\\ x=\frac{1}{2}\]

OpenStudy (anonymous):

oh yes i got it thank

OpenStudy (anonymous):

yw